If p(A) = 0.6 and k is the number of successes of A in n trials


(A) Show that p {550<= k <= 650} = 0.999 , for n = 1000


(B) Find n such that p{0.59n <= k <= 0.61n} = 0.95

1
Expert's answer
2022-02-09T14:54:20-0500

"n=1000\\p=0.6\\q=1-p=1-0.6=0.4"

To solve this question, we shall use the Normal approximation to Binomial.

"\mu=E(x)=np=0.6\times1000=600"

Variance, "\sigma^2=npq=0.6\times0.4\times1000=240"

This implies that the standard deviation "\sigma=\sqrt{\sigma^2}=\sqrt{240}=15.4919334"


"a)"

"p(550\lt k\lt650)=p({550-\mu-0.5\over\sigma}\lt Z\lt{650-\mu+0.5\over\sigma}=p({550-600-0.5\over15.49}\lt Z\lt{650-600+0.5\over15.49})=p(-3.26\lt Z\lt 3.26)=\phi(3.26)-\phi(-3.26)=0.9994-0.0006=0.9988\approx0.999"

as required.

0.5 is the correction factor.


"b)"

"p=0.6\\q=1-p=1-0.6=0.4"

"error=\epsilon={1-0.95\over2}=0.025"

Now,

"p(0.59n \le k \le0.61n) = 2\phi({\epsilon\sqrt{{n\over pq}}})-1=0.95"

This implies that,

"2\phi({\epsilon\sqrt{{n\over pq}}})-1=0.95\implies\phi(0.025\sqrt{n\over0.24})=0.975"

So,

"(0.025\sqrt{n\over0.24})=Z_{0.975}" where "Z_{0.975}" is the table value associated with 0.975. From the tables, "Z_{0.975}=1.96"

Now we have that

"(0.025\sqrt{n\over0.24})=1.96\implies\sqrt{n\over0.24}=78.4\implies n=1475.1744\approx 1475"

For "p(0.59n \leq k \le 0.61n) = 0.95" to hold, the value of "n" must be greater than 1475. That is "n\gt1475".